OpenStudy (mbinck1013):

derivative of 2x(x^2+1)^4dx

OpenStudy (sshayer):

i think it is antiderivative or integral .\[I=\int\limits 2x \left( x^2+1 \right)^4dx\] put \[x^2+1=t,2x~dx=dt\] \[I=\int\limits t^4 dt=\frac{ t^5 }{ 5 }+c\] replace the value of t

OpenStudy (raffle_snaffle):

Face palm.

OpenStudy (learner):

Just in case the \(dx\) is placed accidentally in there, and this is indeed a problem that requires to differentiate the expression, I will give you a few examples. \(\text{___________________________________________________}\) But, before I give examples, note please the following: Suppose you had a function \(\color{black}{\displaystyle y(x)=f(x)\cdot {\rm h}[g(x)]}\) Then, using the "Product Rule", and the "Chain Rule", you will have: \(\color{black}{\displaystyle y'(x)=f'(x)\cdot {\rm h}[g(x)]+f(x)\cdot {\rm h}'[g(x)]\cdot g'(x)}\) \(\text{___________________________________________________}\) Example 1: \(\color{black}{\displaystyle f(x)=ax(x^n+m)^{\beta};~~~\beta\ne-1}\) \(\color{black}{\displaystyle f'(x)=a(x^n+m)^{\beta}+a x\cdot \beta(x^n+m)^{\beta-1}(nx^{n-1})}\) \(\color{black}{\displaystyle f'(x)=a(x^n+m)^{\beta}+a\beta nx^n (x^n+m)^{\beta-1}}\) \(\text{___________________________________________________}\) Example 2: \(\color{black}{\displaystyle f(x)=2x(e^{3x}+x^6)^{4}}\) \(\color{black}{\displaystyle f'(x)=2(e^{3x}+x^6)^4+2x\left[4(e^{3x}+x^6)^{3}\cdot (3e^{3x}+6x^5)\right]}\) \(\color{black}{\displaystyle f'(x)=2(e^{3x}+x^6)^4+24x(e^{3x}+x^6)^{3} (e^{3x}+2x^5)}\)

OpenStudy (sshayer):

if there is no dx then certainly it is a problem of derivatives.